# Vector interaction bounds in NJL-like models from LQCD estimated curvature of the chiral crossover line

### v1

@@ I Introduction
Among the various types of interactions, vector interaction plays a crucial role. Most importantly, in a dense medium, it becomes essential to account for a non-zero vector interaction due to its direct coupling with the number density operator. As a result, even if we start with a zero strength of the vector interaction, the non-zero density environment induces a finite strength for this interaction. As it becomes relevant only at finite density, it cannot be determined using vector meson properties in the vacuum, and its strength remains an uncertain quantity. This contrasts with other interactions, such as the scalar interaction, which can be fixed based on the vacuum properties. Consequently, in much of the existing literature, the strength of the vector interaction ($G_{V}$) is varied in units of the scalar interaction strength $(G_{S})$. It is also worth noting that the sign of $G_{V}$ is not universally agreed upon and can be either positive or negative, leading to repulsive or attractive interactions, respectively [19].
The vector interaction has significant effects on the restoration of chiral symmetry at non-zero baryon chemical potential ($\mu_{B}$) and directly influences the location of the critical endpoint (CEP) in the $T-\mu_{B}$ plane. Furthermore, the curvature coefficients of the chiral crossover line around zero $\mu_{B}$ are strongly affected by the vector interaction. These features have been exploited to understand the role of vector interaction in QCD and to constrain its range in effective model frameworks. For instance, Refs. [20, 21, 22, 23, 24] used different effective model frameworks, including variants of the NJL model, to estimate the vector interaction strength primarily from the LQCD-derived curvature of the chiral crossover line, among other observables. Due to the lack of precise estimates from LQCD, the bounds remain broad but predominantly suggest a repulsive interaction, with the chiral transition becoming a crossover throughout the phase diagram at a sufficiently strong positive interaction strength [19]. Vector interactions also have a profound impact on bulk thermodynamic properties at high densities, contributing to an acceptable equation of state (EoS) for astrophysical objects, such as neutron stars [25] or hybrid stars [26].
@@
We varied the vector interaction strengths $G_{V}$ and $g_{V}$ in units of $G_{S}$ across positive and negative ranges and evaluated the corresponding $\kappa_{2}^{B}$ and $\kappa_{4}^{B}$. Using precise lattice QCD data for $\kappa_{2}^{B}$, we narrowed down the range of these vector interaction strengths. However, $\kappa_{4}^{B}$ estimates from LQCD are consistently zero within uncertainties, offering limited guidance. Lattice results that impose the strangeness neutrality condition (zero net strange quark) provide an additional method for constraining the data with a finite strangeness chemical potential, $\mu_{S}$. This motivates further exploration of the impact of $\mu_{S}$ on the curvature coefficients $\kappa_{2,4}^{B}$ in the presence of vector interactions, previously studied in Ref. [27] without them. This analysis enhances our understanding of flavor mixing between the light and strange quark sectors. In the NJL model, the ’t Hooft determinant term generally captures flavor mixing, while a flavor-independent vector interaction couples the two sectors via their number densities as well. By examining the variation of $\kappa_{2}^{B}$ with $\mu_{S}$, we propose a novel approach to distinguish the contributions from these two mechanisms of flavor mixing and outline potential implications. With the narrowed range of $G_{V}$ and $g_{V}$, we finally provide the location of the critical endpoint (CEP) in our model, which is consistent with the existing LQCD bound [7, 8], however, differing significantly from functional methods’ present predictions [28, 29].
@@ II.1 Details of the model
$-8K[\det(\bar{q}P_{R}q)+\det(\bar{q}P_{L}q)]$ (1)
$-\begin{cases}G_{V}(\bar{q}\gamma^{\mu}q)^{2}\\ g_{V}\sum_{a=0}^{8}\left[(\overline{q}\gamma^{\mu}\lambda_{a}q)^{2}+(\overline{q}i\gamma^{\mu}\gamma^{5}\lambda_{a}q)^{2}\right]\ \end{cases}$
where the quarks are represented by $q$ as $q^{\text{T}}=(u,d,s)$. $\hat{m}=\text{diag}(m_{u},m_{d},m_{s})$ and $\hat{\mu}=\text{diag}(\mu_{u},\mu_{d},\mu_{s})$ are the mass and chemical potential matrices in the flavor space, respectively. $G_{S}$ is the strength of the Lorentz scalar interaction that is invariant under global $U(3)\times U(3)$ symmetry, where $\lambda_{a}$’s are the Gell-Mann matrices, the generator of $U(3)$ symmetry in the fundamental representation with normalization $\text{Tr}[\lambda_{a}\lambda_{b}]=2\delta_{ab}$ and $\lambda_{0}\propto\mathbb{1}_{3\times 3}$. To take into account the effect of axial anomaly, one explicitly breaks the $U(1)_{A}$ symmetry by introducing the Kobayashi–Maskawa–’t Hooft (KMT) interaction [32, 33] with coupling $K$, where $P_{L/R}=(1\mp\gamma_{5})/2$ is the chiral projection operator. We have further considered two types of vector interaction—  flavor independent and flavor dependent [30]. In that regard, they are two different models, and we call them Model-I (flavor independent, $G_{V}$) and Model-II (flavor dependent, $g_{V}$). Next, we obtain the bosonized action by introducing the meanfields (MFs).
@@
[…]
$\left.+\ln\quantity[1+e^{-(E_{i}(p)-\tilde{\mu}_{i})/T}]\right).$ (3)
with $N_{c}=3$, the number of colors; $E_{i}=\sqrt{p^{2}+M_{i}^{2}}$, $\sigma_{i}=\expectationvalue*{\bar{q}_{i}q_{i}}$ and $n_{i}=\expectationvalue*{{q}_{i}^{\dagger}q_{i}}$ are the energy, quark condensate and number density of the $i$-th quark, respectively. To render finite contribution from the vacuum term ($\Omega_{\text{vac}}$), we use $3$-momentum cutoff, $\Lambda$, which also represents the scale of the theory. For $i\neq j\neq k\in\{u,d,s\}$, the $i-$th quark effective mass, $M_{i}$, is given by the gap equation
$M_{i}=m_{i}-4G_{S}\sigma_{i}-2K\sigma_{j}\sigma_{k},$ (4)
@@ III.1 Constraining $G_{V}$ and $g_{V}$ from $\kappa_{2}^{B}$
Let us first consider the curvature coefficients of the crossover line evaluated with the inclusion of vector interactions. This analysis aids in constraining the interaction strengths $G_{V}$ and $g_{V}$ by comparing them with the lattice QCD estimates of $\kappa_{2}^{B}$ and $\kappa_{4}^{B}$, which are available for the crossover line at small values of $\mu_{B}$. In the NJL model, the pseudocritical temperature ($T_{pc}$) for a given set of $\mu_{B}$ and $\mu_{S}$ values is determined as the inflection point of the order parameter, the light quark condensate, $\sigma_{l}$. We extract the curvature coefficients $\kappa_{2}^{B}$ and $\kappa_{4}^{B}$ by parameterizing the respective crossover lines using the ansatz given in Eq. (8), within the range $\mu_{B}/T_{pc}(0)\leq 1.0$.
@@
Let us first summarize the results obtained without the vector interactions in Table 2. We would particularly like to draw attention to $\kappa_{2}^{B,\mu_{S}=0}$ and $\kappa_{2}^{B,n_{s}=0}$. These two curvature coefficients are almost independent of the model parameter sets and have very good agreement with the lattice estimations. On the other hand, $\kappa_{2}^{S}(\mu_{B}=0)$ is sensitive to the choice of model parameters, especially the ’t Hooft determinant, $K$. Due to this robust nature of $\kappa_{2}^{B,\mu_{S}=0}$ and $\kappa_{2}^{B,n_{s}=0}$, we will use them to constrain the strength of the vector interactions and will study the effect of the allowed ranges on $\kappa_{2}^{S,\mu_{B}=0}$. The LQCD data, with which the ranges of $G_{V}$ or $g_{V}$ are constrained, have improved significantly in the last few years. In light of the latest LQCD data, we vary $G_{V}$ and $g_{V}$ in units of $G_{S}$, respectively, and examine the corresponding variations in $\kappa_{2}^{B}$ and $\kappa_{4}^{B}$. In Fig. 1, we illustrate the dependence of $\kappa_{2}^{B}$ on $G_{V}/G_{S}$ and $g_{V}/G_{S}$ for models I and II, respectively.
@@
The points where our estimates intersect with the LQCD bands determine the allowed upper and lower bounds of the vector interaction strengths. The resulting ranges of $G_{V}/G_{S}$ and $g_{V}/G_{S}$ are summarized in Table 3. To understand the effect of model parameters on the estimated range of vector coupling strength, we have used two distinct parameter sets as given in Table 1. From the upper panel of Fig. 1, it is evident that the allowed range of the vector coupling strength is very close to each other for the two parameter sets. It appears that such models exhibit some characteristic features in this regard and are largely independent of the parameters involved [48]. Hence, in the following discussions, we will be presenting all the findings only for parameter set I.
@@
We wish to reiterate that in Model I, there is a noticeable crossing between the cases of $\mu_{S}=0$ and $n_{s}=0$ for both parameter sets, whereas such a feature is absent in Model II, where an almost parallel trend seems to exist. In Ref. [27], the authors demonstrated that a finite value of $\mu_{S}$ tends to decrease $\kappa_{2}^{B}$, which is primarily responsible for the lower value of $\kappa_{2}^{B}$ when considering the $n_{s}=0$ case. According to the analytical expression for the effective chemical potential in Model II, as given in Eq. (7), the strangeness neutrality condition explicitly requires $\mu_{S}=\mu_{B}/3$. This introduces a positive shift in $\mu_{S}$ relative to the $\mu_{S}=0$ case, leading to the observed parallel behavior and the absence of any crossing within the explored range of the coupling strength $g_{V}$. In contrast, for Model I, the $n_{s}=0$ condition does not fix $\mu_{S}$ to a specific value but instead makes it dependent on the light quark number densities as $\mu_{S}=\mu_{B}/3-2G_{V}(n_{u}+n_{d})$. This structure implies that at a certain positive $G_{V}$, $\mu_{S}$ will reach zero, explaining the crossing with the $\mu_{S}=0$ case in Model I.
@@ III.2 $\mu_{S}$ dependence of $\kappa_{2}^{B}$
Figure 2: The curvature coefficient ($\kappa_{2}^{B}$) as a function of strangeness chemical potential ($\mu_{S}$). $[G_{V}^{-},G_{V}^{+}]$ and $[g_{V}^{-},g_{V}^{+}]$ are the allowed ranges of vector interaction found using LQCD data. Red data points are without the vector interaction i.e., $G_{V}=g_{V}=0$. Black point represents the LQCD estimations of the $\kappa^{B}_{2}$ at $\mu_{S}=0$ [4].
In a recent work [27], some of us investigated the dependence of the curvature coefficients on $\mu_{S}$ and proposed a novel approach for determining the strength of the ’t Hooft determinant interaction, $K$. The ’t Hooft term induces mixing between light and heavy quark flavors, making the precise determination of its strength critical for estimating the degree of flavor mixing in a $2+1$ NJL model. On the other hand, in a $2$-flavor NJL model, this effect gets reflected in isospin-sensitive observables only in the isospin asymmetric scenarios, as shown in Ref. [49, 50]. In the presence of a flavor-independent vector interaction, the effective chemical potential, $\tilde{\mu}_{i}$, depends on the densities of all quark flavors. As a result, the term $G_{V}\sum_{i}n_{i}$ introduces a form of flavor mixing along with the ’t Hooft interaction induced mixing through the constituent mass [see Eq. (4)].
@@
It is important to note that the crossing points differ between Model-I and Model-II. In Model-II, this crossing corresponds to a negative value of $\kappa_{2}^{B}$ and occurs at a higher value of $\mu_{S}$. This suggests that $\mu_{S}$ has a weaker influence on the chiral transition line in the presence of flavor-dependent vector interactions $(g_{V}\neq 0)$. Moreover, for $K=0$, the flavor-dependent interaction related to $\mu_{S}$ does not impact the light quark chemical potential, leading to a curvature coefficient that remains independent of $\mu_{S}$. This is demonstrated by the two straight lines, representing $g_{V}^{+}$ (soft blue) and $g_{V}^{-}$ (soft orange), in Fig. 3. This indicates that in Model-II, there is no flavor mixing due to nonzero $g_{V}$ and the reduction in $\kappa_{2}^{B}$ with increasing $\mu_{S}$, as presented in the lower panel Fig.2, is entirely due to the flavor mixing term $K$.

### v2

@@ I Introduction
Among the various types of interactions, vector interaction plays a crucial role. Its strength can be generally fixed from the vector mesons as they emerge as peaks in the quark–antiquark spectral functions [19, 12]. But the masses of these vector mesons are less robustly constrained than those of pseudoscalar mesons in an effective model scenario, since they typically lie close to the model cutoff [20]. Along with this, the vector meson modes in the NJL model are embedded in an unphysical quark–antiquark continuum arising from the lack of confinement. Consequently, the lowest-lying vector mesons, such as the $\rho$, can decay into this continuum. As a result, the model does not reliably reproduce the vector meson mass spectrum as it does for the pseudoscalar sector [19]. Although in these studies, the extracted vector coupling had a value $G_{V}\sim(1.5)\,G_{S}$ depending on the chosen parameter set. This provides one of the earliest quantitative determinations of the interaction strength [20, 19].
Most importantly, in a dense medium, it becomes essential to account for a non-zero vector interaction due to its direct coupling with the number density operator, $(\bar{\psi}\gamma_{0}\psi)$. As a result, even if we start with a zero strength of the vector interaction, the non-zero density environment induces a finite strength for this interaction. As it becomes relevant only at finite density, this induced vector interaction cannot be determined using vector meson properties in the vacuum, and its strength remains an uncertain quantity [21, 22]. Ref. [23] introduces a separate isovector-vector interaction to differentiate between the isoscalar-vector $(\omega)$ and isovector-vector $(\rho)$ channels. However, the strength of these two interactions are treated as a free parameter and varied as a ratio to the scalar coupling constant. Consequently, in much of the existing literature, the strength of the vector interaction ($G_{V}$) is varied in units of the scalar interaction strength $(G_{S})$ [24, 25, 26, 27, 28, 29, 30]. It is also worth noting that the sign of the induced vector interaction is not universally agreed upon and can be either positive or negative, leading to repulsive or attractive interactions, respectively, as elaborated well in Ref. [21].
The vector interaction has significant effects on the restoration of chiral symmetry at non-zero baryon chemical potential ($\mu_{B}$) and directly influences the location of the critical endpoint (CEP) in the $T-\mu_{B}$ plane. Furthermore, the curvature coefficients of the chiral crossover line around zero $\mu_{B}$ are strongly affected by the vector interaction. These features, particularly the curvature coefficients, have been exploited to understand the role of vector interaction in QCD and to constrain its range in effective model frameworks, and they form a major part of the present study. Refs. [31, 24, 25, 26, 28, 27] used different effective model frameworks, including variants of the NJL model, to estimate the vector interaction strength primarily from the LQCD-derived curvature of the chiral crossover line, among other observables. Due to the lack of precise estimates from LQCD, the bounds from the curvatue coefficients remain broad but predominantly suggest a repulsive interaction, with the chiral transition becoming a crossover throughout the phase diagram at a sufficiently strong positive interaction strength [21]. Vector interactions also have a profound impact on bulk thermodynamic properties at high densities, contributing to an acceptable equation of state (EoS) for astrophysical objects, such as neutron stars [32] or hybrid stars [33, 34].
@@
We varied the vector interaction strengths $G_{V}$ and $g_{V}$ in units of $G_{S}$ across positive and negative ranges and evaluated the corresponding $\kappa_{2}^{B}$ and $\kappa_{4}^{B}$. Using precise lattice QCD data for $\kappa_{2}^{B}$, we narrowed down the range of these vector interaction strengths. However, $\kappa_{4}^{B}$ estimates from LQCD are consistently zero within uncertainties, offering limited guidance. Lattice results that impose the strangeness neutrality condition (zero net strange quark) provide an additional method for constraining the data with a finite strangeness chemical potential, $\mu_{S}$. This motivates further exploration of the impact of $\mu_{S}$ on the curvature coefficients $\kappa_{2,4}^{B}$ in the presence of vector interactions, previously studied in Ref. [35] without them. This analysis enhances our understanding of flavor mixing between the light and strange quark sectors. In the NJL model, the ’t Hooft determinant term generally captures flavor mixing, while a flavor-independent vector interaction couples the two sectors via their number densities as well. By examining the variation of $\kappa_{2}^{B}$ with $\mu_{S}$, we propose a novel approach to distinguish the contributions from these two mechanisms of flavor mixing and outline potential implications. With the narrowed range of $G_{V}$ and $g_{V}$, we finally determine the location of the critical endpoint (CEP) in our model, which is consistent with the existing LQCD bound [7, 8], but differs from the current predictions of functional methods [36, 37].
@@ II.1 Details of the model
$-8K\left[\det(\bar{q}P_{R}q)+\det(\bar{q}P_{L}q)\right]$
$-\left\{\begin{array}[]{ll}G_{V}(\bar{q}\gamma^{\mu}q)^{2}\\[4.30554pt] g_{V}\sum_{a=0}^{8}\left[(\bar{q}\gamma^{\mu}\lambda_{a}q)^{2}+(\bar{q}i\gamma^{\mu}\gamma^{5}\lambda_{a}q)^{2}\right]\end{array}\right$. (3)
where the quarks are represented by $q$ as $q^{\text{T}}=(u,d,s)$. $\hat{m}=\text{diag}(m_{u},m_{d},m_{s})$ and $\hat{\mu}=\text{diag}(\mu_{u},\mu_{d},\mu_{s})$ are the mass and chemical potential matrices in the flavor space, respectively. $G_{S}$ is the strength of the Lorentz scalar interaction that is invariant under global $U(3)\times U(3)$ symmetry, where $\lambda_{a}$’s are the Gell-Mann matrices, the generator of $U(3)$ symmetry in the fundamental representation with normalization $\text{Tr}[\lambda_{a}\lambda_{b}]=2\delta_{ab}$ and $\lambda_{0}\propto\mathbb{1}_{3\times 3}$. To take into account the effect of axial anomaly, one explicitly breaks the $U(1)_{A}$ symmetry by introducing the Kobayashi–Maskawa–’t Hooft (KMT) interaction [40, 41] with coupling $K$, where $P_{L/R}=(1\mp\gamma_{5})/2$ is the chiral projection operator. We have further considered two types of vector interaction—  flavor independent and flavor dependent [38]. In that regard, they are two different models, and we call them Model-I (flavor independent, $G_{V}$) and Model-II (flavor dependent, $g_{V}$). Next, we utilize the meanfield approximation to substitute the auxiliary bosonic fields, obtained through the Hubbard-Stratonovich transformation of the Lagrangian in Eq. (3), with the corresponding expectation values.
@@
$\Omega_{\text{med}}$ $=-2N_{c}T\sum_{i}\int_{0}^{\infty}\frac{d^{3}p}{(2\pi)^{3}}\left(\ln\left[1+e^{-(E_{i}(p)+\tilde{\mu}_{i})/T}\right]\right$.
$\left.+\ln\left[1+e^{-(E_{i}(p)-\tilde{\mu}_{i})/T}\right]\right),$ (5)
with $N_{c}=3$, the number of colors; $E_{i}=\sqrt{p^{2}+M_{i}^{2}}$, $\sigma_{i}=\langle\bar{q}_{i}q_{i}\rangle$ and $n_{i}=\langle{q}_{i}^{\dagger}q_{i}\rangle$ are the energy, quark condensate and number density of the $i$-th quark, respectively. To render finite contribution from the vacuum term ($\Omega_{\text{vac}}$), we use $3$-momentum cutoff, $\Lambda$, which also represents the scale of the theory. For $i\neq j\neq k\in\{u,d,s\}$, the $i$-th quark effective mass, $M_{i}$, is given by the gap equation
$M_{i}=m_{i}-4G_{S}\sigma_{i}+2K\sigma_{j}\sigma_{k},$ (6)
@@ III.1 Constraining $G_{V}$ and $g_{V}$ from $\kappa_{2}^{B}$
Let us first consider the curvature coefficients of the crossover line evaluated with the inclusion of vector interactions. This analysis aids in constraining the interaction strengths $G_{V}$ and $g_{V}$ by comparing them with the lattice QCD estimates of $\kappa_{2}^{B}$ and $\kappa_{4}^{B}$, which are available for the crossover line at small values of $\mu_{B}$. In the NJL model, the pseudocritical temperature ($T_{pc}$) for a given set of $\mu_{B}$ and $\mu_{S}$ values is determined as the inflection point of the order parameter, the light quark condensate, $\sigma_{l}$ [^1] It is to be noted that Ref. [4] uses a combination of light and strange quark condensates to define the order parameter. This is done to get rid of the divergences that arise in lattice QCD calculation [71].. We extract the curvature coefficients $\kappa_{2}^{B}$ and $\kappa_{4}^{B}$ by parameterizing the respective crossover lines using the ansatz given in Eq. (10), within the range $\mu_{B}/T_{pc}(0)\leq 1.0$.
@@
We, first, summarize the results obtained without the vector interactions in Table 2. We would particularly like to draw attention to $\kappa_{2}^{B,\mu_{S}=0}$ and $\kappa_{2}^{B,n_{s}=0}$. These two curvature coefficients are almost independent of the model parameter sets and have very good agreement with the lattice estimations. On the other hand, $\kappa_{2}^{S}(\mu_{B}=0)$ is sensitive to the choice of model parameters, especially the ’t Hooft determinant, $K$. Due to this robust nature of $\kappa_{2}^{B,\mu_{S}=0}$ and $\kappa_{2}^{B,n_{s}=0}$, we will use them to constrain the strength of the vector interactions and will simply study the effect of these interactions on $\kappa_{2}^{S,\mu_{B}=0}$. The LQCD data, with which the ranges of $G_{V}$ or $g_{V}$ are constrained, have improved significantly in the last few years [57, 45, 5, 4, 6]. In light of the latest LQCD data, we vary $G_{V}$ and $g_{V}$ in units of $G_{S}$, respectively, and examine the corresponding variations in $\kappa_{2}^{B}$ and $\kappa_{4}^{B}$. In Fig. 1, we illustrate the dependence of $\kappa_{2}^{B}$ on $G_{V}/G_{S}$ and $g_{V}/G_{S}$ for Model-I and Model-II, respectively. Compared to Refs. [31, 24, 25, 26, 28, 27], which consider only flavor independent vector interaction, we explore both the flavor-independent and flavor-dependent interactions. We also do not include the background gauge field, intending to investigate the effects of the vector interactions on the chiral crossover line solely using the chiral dynamics, particularly because the NJL model is known to capture these dynamics reliably.
@@
The points where our estimates intersect with the LQCD bands determine the allowed upper and lower bounds of the vector interaction strengths. The resulting ranges of $G_{V}/G_{S}$ and $g_{V}/G_{S}$ are summarized in Table 3. To understand the effect of model parameters on the estimated range of vector coupling strength, we have used two distinct parameter sets as given in Table 1. From the upper panel of Fig. 1, it is evident that the allowed ranges of the vector coupling strength for the two parameter sets are very close to each other, while being constrained by the lattice QCD estimation of $\kappa_{2}^{B}$. It appears that such models exhibit some characteristic features in this regard and the values of $\kappa_{2}^{B}$ are largely independent of the parameters involved [58]. Hence, in the following discussions, we will be presenting all the findings only for parameter set I.
We find the allowed ranges of the vector interactions accommodate both attractive and repulsive interactions for both models. The estimated range of $G_{V}$ for Model-I is in line with the findings of Refs. [31, 24, 25, 26, 28, 27], however, most of which favored a stronger repulsive vector interaction as compared to the present results. This is predominantly due to the available LQCD data which have evolved significantly over time and settled for comparatively higher values [57, 45, 5, 4, 6].
@@
We wish to reiterate that in Model-I, there is a noticeable crossing between the cases of $\mu_{S}=0$ and $n_{s}=0$ for both parameter sets, whereas such a feature is absent in Model-II, where an almost parallel trend seems to exist. In Ref. [35], the authors demonstrated that a finite value of $\mu_{S}$ tends to decrease $\kappa_{2}^{B}$, which is primarily responsible for the lower value of $\kappa_{2}^{B}$ when considering the $n_{s}=0$ case. According to the analytical expression for the effective chemical potential in Model-II, as given in Eq. (9), the strangeness neutrality condition explicitly requires $\mu_{S}=\mu_{B}/3$. This introduces a positive shift in $\mu_{S}$ relative to the $\mu_{S}=0$ case, leading to the observed parallel behavior and the absence of any crossing within the explored range of the coupling strength $g_{V}$. In contrast, for Model-I, the $n_{s}=0$ condition does not fix $\mu_{S}$ to a specific value but instead makes it dependent on the light quark number densities as $\mu_{S}=\mu_{B}/3-2G_{V}(n_{u}+n_{d})$. This structure implies that at a certain positive $G_{V}$, $\mu_{S}$ will reach zero, explaining the crossing with the $\mu_{S}=0$ case in Model-I.
@@
At this point, it is worth noting that a similar variation of $\kappa_{2}^{B,\mu_{S}=0}$ with varying strength of vector interaction (as shown in Fig.1) had earlier been explored in Polyakov loop extended NJL (PNJL) model  [24, 25, 26, 27]. These studies further employed the lattice QCD estimation of $\kappa_{2}^{B,\mu_{S}=0}$, available at the time, to constrain the range of those vector interactions. Our results exhibit a similar behavior and are in line with those from PNJL. Although the constraints values $G_{V}$ and $g_{V}$ from these PNJL studies favored a positive range. This is primarily because the lattice results available at that time were significantly lower in values than the present estimates.
@@ III.2 $\mu_{S}$ dependence of $\kappa_{2}^{B}$
Figure 2: The curvature coefficient ($\kappa_{2}^{B}$) as a function of strangeness chemical potential ($\mu_{S}$). $G_{V}^{-},G_{V}^{+}]$ and $g_{V}^{-},g_{V}^{+}]$ are the allowed ranges of vector interaction found using LQCD data. Red data points are without the vector interaction i.e., $G_{V}=g_{V}=0$. Black point represents the LQCD estimations of the $\kappa^{B}_{2}$ at $\mu_{S}=0$ [4].
In a recent work [35], some of us investigated the dependence of the curvature coefficients on $\mu_{S}$ and proposed a novel approach for determining the strength of the ’t Hooft determinant interaction, $K$. The ’t Hooft term induces mixing between light and heavy quark flavors, making the precise determination of its strength critical for estimating the degree of flavor mixing in a $2+1$ NJL model. On the other hand, in a $2$-flavor NJL model, such mixing effect between the two light flavors gets reflected only in the isospin asymmetric scenarios, as shown in Ref. [59, 60]. In the presence of a flavor-independent vector interaction, the effective chemical potential, $\tilde{\mu}_{i}$, depends on the densities of all quark flavors. As a result, the term $G_{V}\sum_{i}n_{i}$ introduces a form of flavor mixing along with the ’t Hooft interaction induced mixing through the constituent mass [see Eq. (6)].
@@
It is important to note that the crossing points differ between Model-I and Model-II. In Model-II, this crossing corresponds to a negative value of $\kappa_{2}^{B}$ and occurs at a higher value of $\mu_{S}$ (Fig. 2). This suggests that $\mu_{S}$ has a weaker influence on the chiral transition line in the presence of flavor-dependent vector interactions $(g_{V}\neq 0)$. Moreover, for $K=0$, the flavor-dependent interaction related to $\mu_{S}$ does not impact the light quark chemical potential, leading to a curvature coefficient that remains independent of $\mu_{S}$. This is demonstrated by the two straight lines, representing $g_{V}^{+}$ (soft blue) and $g_{V}^{-}$ (soft orange), in Fig. 3. This indicates that in Model-II, there is no flavor mixing due to nonzero $g_{V}$ and the reduction in $\kappa_{2}^{B}$ with increasing $\mu_{S}$, as presented in the lower panel Fig. 2, is entirely due to the flavor mixing term $K$.
We note that a recent lattice QCD study [61] has investigated the chiral phase transition as a function of both $\mu_{B}$ and $\mu_{S}$, and determined the leading-order curvature coefficients. From their parametric representation of the phase line, one can extract $\kappa_{2}^{B}$ for chosen values of $\mu_{S}$. Interestingly, this parametrization exhibits the same decreasing trend of $\kappa_{2}^{B}$ with increasing $\mu_{S}$, with $\kappa_{2}^{B}$ turning zero around $\mu_{S}\simeq 120~\text{MeV}$. The estimations of higher-order curvature coefficients from lattice QCD at larger $\mu_{S}$ would allow a more quantitative assessment of the strength of the ’t Hooft and vector interactions.